OpenStudy (gabylovesyou):

Find the factors of 3a2 + 14a − 24.

OpenStudy (goformit100):

Madam Can I solve it for you ?

OpenStudy (goformit100):

Madam See Q = 3a^2 - 14a - 24 3Q = 9a^2 - 14(3a) - 72 3Q = ( 3a)^2 - 14(3a) - 72 Let u = 3a 3Q = u^2 - 14u - 72 3Q = ( u - 18 )(u + 4 ) 3Q = ( 3a - 18 )(3a + 4 ) 3Q = 3(a - 6 )(3a + 4 ) Q = ( a - 6 )(3a + 4 )

OpenStudy (goformit100):

@Gabylovesyou there is Another way of doing this is.... wanna know ?

OpenStudy (gabylovesyou):

thats not one of my choices.... yes

OpenStudy (goformit100):

[ ( -14 / 2 ) - k ] [ ( - 14 / 2 ) + k ] = 3( - 24 ) ( - 7 - k ) ( - 7 + k ) = - 72 49 - k^2 = - 72 121 = k^2 11 = k - 7 - 11 = - 18 - 7 + 11 = 4 3a^2 - 18a + 4a - 24 = 3a( a- 6 ) + 4 ( a - 6 ) = ( 3a + 4 ) ( a - 6 )

OpenStudy (goformit100):

Is it ok now ?

OpenStudy (gabylovesyou):

no... lol

OpenStudy (goformit100):

Now last option, @nader1 My friend please Help Gaby Madam.

OpenStudy (anonymous):

ok wait let me read everything

OpenStudy (anonymous):

3a^2 + 14a − 24 3a^2 + 18a - 4a - 24 (3a - 4) (a + 6)

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

tell me :) gaby

OpenStudy (goformit100):

On Behalf of Gaby Madam Thankyou @nader1

OpenStudy (anonymous):

anytime i wait to see if i am correct but i am sure

jimthompson5910 (jim_thompson5910):

3a^2+14a-24 3a^2+18a-4a-24 (3a^2+18a)+(-4a-24) 3a(a+6)+(-4a-24) 3a(a+6)-4(a+6) (3a-4)(a+6) --------------------------------------------------------------------------- 3a^2+14a-24 completely factors to (3a-4)(a+6)